Write the equation of the line passing through (3,-7),(-6,-7)
Answer by checkley71(8403)

Y=mX+b the slope is (-7,-7)/(-6,3)=0/-3 therefore the slope is 0 and the y intercept is -7
thus the equation is Y=0x-7 or y=-7